OpenStudy (anonymous):

Which is not a common factor for 150 and 450?

OpenStudy (anonymous):

what are the options

OpenStudy (whpalmer4):

109736 is one correct answer :-)

OpenStudy (anonymous):

XD :)

OpenStudy (anonymous):

9

OpenStudy (whpalmer4):

Another good answer!

OpenStudy (anonymous):

lol thx ^_^

OpenStudy (whpalmer4):

Without the answer options, how about we factor both numbers? Then someone in possession of the answer options can pick the one which does not appear in our list of factors! 150 = 2*75 = 2*3*25 = 2*3*5*5 450 = 2*225 = 2*3*75 = 2*3*3*25 = 2*3*3*5*5 Any number you can't make up by multiplying the common factors in that list (2, 3, 5, 5) is not a common factor of 150 and 450. For example, you can't make up 9 because there is only 1 3 in common. 15 is a possible factor, because you can multiply 3*5 =15. Etc.

OpenStudy (anonymous):

im likin this dude :P

OpenStudy (anonymous):

(thats so not supposed to sound creepy)

OpenStudy (anonymous):

well it did

OpenStudy (anonymous):

XD ikr :P

OpenStudy (anonymous):

hahahaha :D

OpenStudy (whpalmer4):

Creepiness is in the ear of the beholder, isn't it? I thought it was fine :-)